OpenStudy (anonymous):

Please explain how to do a problem like i÷6=2, I have to show my work. This is a 6th grade math problem.

OpenStudy (anonymous):

i / 6 = 2 you need to get i on its own - to the left of '=' to do this you multiply both sides of = sign by 6 so i/6 * 6 = 2 * 6 i = 12

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

no probs - plz close the question by clicking on good answer

OpenStudy (anonymous):

how?